OpenStudy (anonymous):

In a random sample of 678 adults, it was determined that 58 of them have hypertension. Construct a 95% confidence interval. I'm not quite sure how to get xbar for use in the formula, and I don't know what t a/2 refers to.

OpenStudy (anonymous):

i'm not 100% sure how to do this but i'd imagine you use the formula for the standard deviation of the binomial distribution (look on wikipedia) and then do the same thing as the previous problem

OpenStudy (anonymous):

looks like the variance is np(1 − p) so the std dev is sqrt(np(1 − p))

OpenStudy (anonymous):

n=678, p=58/678

OpenStudy (anonymous):

Alright I'll try that thanks

OpenStudy (anonymous):

do you have the answer?

OpenStudy (anonymous):

I can look in the back of the book.

OpenStudy (anonymous):

hold on let me guess it first

OpenStudy (anonymous):

plus or minus 14.3 i think? so 43.7 to 72.3

OpenStudy (anonymous):

According to this it's .065 to .107.

OpenStudy (anonymous):

oh, they want percentages... hold on let me convert

OpenStudy (anonymous):

BANG! that's what i got, 0.0644 to 0.1066

OpenStudy (anonymous):

Awesome :D Thanks.

OpenStudy (anonymous):

So, uhh, what is variance? Does that have to do with degrees of freedom? And standard deviation is just the square root of variance?

OpenStudy (anonymous):

standard deviation is just the sq root of the variance yes

OpenStudy (anonymous):

Okay, so I got the variance and standard deviation and everything, how can I get xbar for use in xbar +- z * stdev/sqrt n?

OpenStudy (anonymous):

actually in this case the formula for std dev already has the sample size built into it. so you just do 1.96 * stddev (plus or minus). where std dev is sqrt(n p (1-p))

OpenStudy (anonymous):

xbar is just 58/678

OpenStudy (anonymous):

Isn't xbar the mean? How can it be so low?

OpenStudy (anonymous):

And where did you get 1.96?

OpenStudy (anonymous):

xbar is the mean probability of finding someone with hypertension

OpenStudy (anonymous):

1.96 is the number of std devs up and down to get to 95% confidence; it's in your quick reference chart

OpenStudy (anonymous):

Ah, okay.

OpenStudy (anonymous):

Uhh, okay, so I have xbar = .085 +- 1.96 * (7.26/sqrt20) and I get .085 +- 3.175, that's not right...

OpenStudy (anonymous):

n = 678 p = 58/678 variance = np (1-p) = 17980/339 = 53.04 stddev = sqrt(variance) = 7.28 95% conf interval = 1.95 * stddev = 14.27 * 1.96 = 14.27 so 95% conf is 58+14.27 (72.27) to 58-14.27 (43.72) converting that back to probabilities you get 72.27/678 = 0.106 to 43.72/678 = 0.0644

OpenStudy (anonymous):

O_o Sure, let me think.

OpenStudy (anonymous):

Where did the 1.95 come from? All I see in the formula is 1.96 for Z.

OpenStudy (anonymous):

typo it should have said 1.96

OpenStudy (anonymous):

Oh, right.

OpenStudy (anonymous):

And I thought xbar was 58/678?

OpenStudy (anonymous):

it is

OpenStudy (anonymous):

But you said 58 up there, not .085.

OpenStudy (anonymous):

the mean is 58/678 yes. but for this particular problem i used the formula for the binomial distribution, which has built into it the number of samples. the std dev of the binomial distribution tells you how many successes you can expect with that particular number of samples (678). the mean is 58, the std dev ends up being 7. in the end you can convert that back to the same units as the sample mean by dividing the number of samples back out.

OpenStudy (anonymous):

Oh, okay... I think I'm going to have a hard time remebering this.

OpenStudy (anonymous):

there's probably some other way to do it in the units of percentages all the way through, i just don't happen to know it

OpenStudy (anonymous):

Well, I got the answer, that's what matters. Thanks.

OpenStudy (anonymous):

My god it's almost midnight and I've done half of these problems.

OpenStudy (anonymous):

I'm gonna go get some food real quick.